Calculate  \oint_{C}^{.} \overrightarrow{F}\cdot d\overrightarrow{x}, where C is the curve below :

F* (22.3 + 2.ru2-20,)ịt (20/3 + 2.r2 ч + 2.r)j
and the area inside C is A (unknown).


F* (22.3 + 2.ru2-20,)ịt (20/3 + 2.r2 ч + 2.r)j
Answer #1

We can use the stoke's theorem to calculate the line integral

F.di

\large \\(\nabla \times \vec{F})= \begin{bmatrix} i &j &k \\ \frac{\partial}{\partial x} & \frac{\partial}{\partial y}&\frac{\partial}{\partial z} \\ 2x^3+2xy^2-2y & 2y^3+2x^2y+2x & 0 \end{bmatrix} \\ = i(0-0)-j(0-0)+k(4xy+2-4xy+2)=4 k

V x F).ndA-/ 4k.kdA 4A

F.di 4A